3. (14 total points) Suppose that Lynn enjoys coconut oil in her coffee. She has very particular preferences, and she must have exactly three spoonfuls of coconut oil for each cup of coffee. Let C be the number of cups of coffee, and O be the number of spoonfuls of coconut oil. Also, let Pc be the price of a cup of coffee. Suppose Lynn has S12 to spend on coffee and coconut oil. Also, the price of coconut oil is S.20 per spoonful. Graph Lynns Price consumption curve for prices, Pc- S.60, Pc-S.90, and Pc-$1.80. Please put the number of cups of coffee (C) on the horizontal axis, and the number of spoonfuls of coconut oil (O) on the vertical axis. Be sure to label your graph carefully and accurately
Answer #1

Let

No. of cups of coffee=C

No. of spoonful of coconut oil=O

Case-1

Price of a cup of coffee=Pc=$0.60 per cup

Price of a spoonful of coconut oil=Po=$0.20 per spoonful

Amount available to spend=I=$12

She needs exactly 3 spoonful of oil for each cup of coffee.

Her consumption line will be given by

C*Pc+3*O*Pc=I2

0.60C+3*O*0.20=I2

0.6C+0.6O=I2

Her consumption line at Pc=$0.60 will be given by

0.6C+0.6O=I2

if C=0, O=20

If O=0, C=20

Case 2

Price of a cup of coffee=Pc=$0.90 per cup

Her consumption line will be given by

C*Pc+3*O*Pc=I2

0.90C+0.6Q=12

Her consumption line at Pc=$0.90 will be given by

0.9C+0.6O=I2

if C=0, O=20

If O=0, C=13.3333

Case 3

Price of a cup of coffee=Pc=$1.80 per cup

Her consumption line will be given by

C*Pc+3*O*Pc=I2

1.80C+0.6O=12

Her consumption line at Pc=$1.80 will be given by

1.8C+0.6O=I2

if C=0, O=20

If O=0, C=6.666667

gv9dkMIABe4A9wB5gD7AH2APsAfZAqntg0ZsT25z